Question 10

Which of the following is CORRECT? When discounting an amount to be received in one year's time at a rate that is quoted as 12% compounding quarterly, we can:

Select one:

A. Discount the amount using an effective monthly rate of 1% where the number of periods is 12.

B. Discount the amount using an effective annual rate (EAR) of (1+0.03)^4-1 =12.5509% where number of periods is 1.

C. Discount the amount using the annual rate of 12% where number of periods is 1.

D. None of the above
Answer #1

CORRECT OPTION : B. Discount the amount using an effective annual rate (EAR) of (1+0.03)^4-1 =12.5509% where number of periods is 1.

As amount is received at the end of the year and rate is compounded quarterly, we have to find effective annual rate

so we have to use : EAR = (1+(r/m))^m - 1 = (1+(0.12/4))^4 - 1 = 12.5509%

Answer : B : [[[[Thumbs up please]]]